# TeViR: Text-to-Video Reward with Diffusion Models for Efficient Reinforcement Learning

### v1

@@ VI Conclusion
Future work may explore several exciting directions to further enhance TeViR. One area of focus could be training video generation models with internet-scale data for a broader range of tasks, which enables TeViR to tackle more complex and realistic operation tasks, and drives advancements in the field of robotics by allowing for more sophisticated and real-world robotic applications. Another promising direction is combining UniPi and TeViR to enhance sample efficiency and training robustness, enabling effective policy learning with a few samples. For instance, UniPi can be used to collect samples and warm up the policy, followed by TeViR for fine-tuning. This aims to achieve more efficient utilization of available samples and foster quicker adaptation to new tasks, which is crucial for advancing the capabilities of robotic systems in dynamic environments under resource-constrained condition. These enhancements aim to push the boundaries of what is achievable in reinforcement learning and robotic manipulation, driving forward the field towards more sophisticated and capable AI systems.

### v2

@@ VI Conclusion
Future work may explore several exciting directions to further enhance TeViR. One area of focus could be training video generation models with internet-scale data for a broader range of tasks, which enables TeViR to tackle more complex and realistic operation tasks, and drives advancements in the field of robotics by allowing for more sophisticated and real-world robotic applications. A more effective key frame selection or prediction mechanisms that are task-aware and semantically grounded is also important, to improve the fidelity and relevance of the generated videos and further reduce ambiguity during reward computation. Another promising direction is combining UniPi and TeViR to enhance sample efficiency and training robustness, enabling effective policy learning with a few samples. For instance, UniPi can be used to collect samples and warm up the policy, followed by TeViR for fine-tuning. This aims to achieve more efficient utilization of available samples and foster quicker adaptation to new tasks, which is crucial for advancing the capabilities of robotic systems in dynamic environments under resource-constrained condition. These enhancements aim to push the boundaries of what is achievable in reinforcement learning and robotic manipulation, driving forward the field towards more sophisticated and capable AI systems.